Left Direct Gregorian Technical Observing with IRTC: UT 2010 March 02

Observer: DMiller JHill (Tucson)
Telescope Operator: WWack
Instrument Support:
Telescope Support:
SW support: MDeLaPena, NCushing, CBiddick, TEdgin on call


Didn't open because of humidity until 4AM MST. Once we opened, we had big problems with GCS. First we had sextractor problems, then there were problems with timeouts waiting for the guider image. 12:00 UT Then the azimuth axis got stuck, so we surrender.


All times in UT

Closed-dome tests of Build BP1

01:00 Michele, Norm, Chris and Doug are in Tucson

01:10 Dave A and James H are working on the Elevation HBS to get the telescope moving

01:10 100% Humidity so could not open, even if telescope was movable

01:10 The LBT Switcher was tested and found to have a problem. Paul fixed it.

02:00 Olga K arrives

02:10 Found that the IIF GUI dies with a core file when MCSPU is restarted. Wayne restarted MCSPU several times and most IIF Gui's died all of the time. Chris found that MCSPU/MCS was putting a bad El number in the DD when MCSPU was restarted. He changed the IIF GUI to capture this problem and sent a patch to the mountain. We tested it and it works fine.

03:00 John H arrives

04:00 Talk to Dave A on the phone and he said the El was still having problems. It could be quick to fix, or could be long.

04:10 Software people change the build on the mountain to AO2 and then go home. Now only Doug and John in Tucson. We see telescope moving up and down, but no news yet. Still 100% humidity and clouds.

Testing IDL scripts while waiting on humidity

06:40 Manually cycled power on AGW4 to revive stuck AzCam. We don't know what made AzCam get stuck in the first place.

06:45 After 2 hours of continuous defrosting, the SX mirror ventilation is deiced and working OK again. See Issue #2474.

06:50 DAshby and friends have telescope working again. Wayne is filling LBC-Red. But humidity is still 100%.

07:30 Wayne and Dave reset the LDG rotator (was against a limit). Wayne cycles the power on the IRTC. It now works

08:54 Doug tries his new IDL presetTelescope routines. Failed because the target.dat file was not being closed (thus file was empty). Failed because the format of the target.dat file was incorrect (must have only one space between entries)

08:59 presetTelesope with IRAF. Successful. Use this to compare my IDL steps and output for debugging.

09:22 IDL presetTelescope was successful (as IRTC@LDG)!

09:34 Registered my IDL routines as Right prime focus, LBC and sent a successful preset to the same star

09:38 Found that the target.dat file being created didn't have large enough floating point values in the FORMAT. Changed to writing string. Fixed!

09:43 Sent preset to new star (as RIGHT LBC @ prime): successful

09:47 Send preset to different star (as LEFT IRTC): right side dropped to none mode. We are surprised that the dropping to none mode was not symmetrical.

10:00 Humidity at 96%.

10:15 John preset the telescope with LBC. Build AO2 successfully pointed the telescope


11:00 Open to partly cloudy sky, but humidity of merely 84%. Right shutter door wouldn't open. T=-4.5 degC.

11:05 pointto M5_1396 ACQUIRE - star is off the field, but nearby. Adjust IE=0 CA=7 Initial collimation was OK.

11:12 pointto M5_1399 ACQUIRE PAM- EL=25 IA=-375 (probably didn't need adjustment as it had been -404)

11:17 pointto BS9158 ACQUIRE PAM- IE=-7 CA=1

Problem of not identifying star on acquisition image

11:24 pointto BS9158 ACTIVE - guiding wasn't stable?

11:26 pointto BS9158 ACTIVE - failed to request full guidecam readout ??

11:28 pointto BS9158 ACTIVE - star was not in the guide box even though it is clearly on the acq image. Wayne correctly says "so how does this happen?".

11:30 pointto BS9158 ACQUIRE (tweaked IE/CA to make it centered precisely)

11:32 pointto BS9158 ACTIVE - the offset fails miserably . GCS/sextractor fails to find the star by several hundred pixels. Fortunately Torsten tells us what to do in the release notes for Build AO2.

11:42 GCSGUI dies while I was editting lbt.conf to change GCSL.sextractorClass from 0.1 to 0.0.

11:43 pointto BS9158 ACTIVE - failed to request full guidecam readout

11:45 pointto BS9158 ACTIVE - succeeds with parameter in lbt.conf set to 0.0, and the seeing has improved too.

Problems getting guider images

11:47 the downtown copy of GCSGUI is stuck.

11:51 GCS is still getting WFS images, but has stopped guiding - that is no guider images are appearing in Repository for 5 minutes. This became Issue #2543 the next day.
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:50 guiderimage000108.fits
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:50 guiderimage000109.fits
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:50 guiderimage000110.fits
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:50 guiderimage000111.fits
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:50 guiderimage000112.fits
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:50 guiderimage000113.fits
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:50 guiderimage000114.fits
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:50 guiderimage000115.fits
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:51 guiderimage000116.fits
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:51 guiderimage000117.fits
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:51 guiderimage000118.fits
-rw-rw-r-- 1 telescope lbt        48960 Mar  2 11:51 guiderimage000119.fits
-rw-rw-r-- 1 telescope lbt       570240 Mar  2 11:51 wfscimage000007.fits
-rw-rw-r-- 1 telescope lbt       570240 Mar  2 11:52 wfscimage000008.fits
-rw-rw-r-- 1 telescope lbt       570240 Mar  2 11:53 wfscimage000009.fits
-rw-rw-r-- 1 telescope lbt       570240 Mar  2 11:54 wfscimage000010.fits
-rw-rw-r-- 1 telescope lbt       570240 Mar  2 11:54 wfscimage000011.fits
-rw-rw-r-- 1 telescope lbt       570240 Mar  2 11:55 wfscimage000012.fits
-rw-rw-r-- 1 telescope lbt       570240 Mar  2 11:56 wfscimage000013.fits
11:58 kill and restart GCSL

11:59 pointto BS9158 ACTIVE

Guiding started again, but GCS is putting up lots of messages about timed out waiting for images.

12:03 kill GCSL and stopAGW -u 4, then restart. "No AGW unit selected". Reauthorized IRTC. (We're confused about what is going on here for this little side problem.) Did it fail to home?

12:08 Authorize LUCIFER, then back to IRTC LDG. This causes GCS to home AGW4.

12:11 Dropping GCSL ????

Problem with Azimuth Axis

12:12 pointto BS9158 ACTIVE - but preset doesn't return. Azimuth Axis is stuck


12:20 We surrender and close the dome.

12:20 readGuideCam left -e 1000 100 gives "guide image invalid" although I can take them one at a time.

-- DougMiller - 02 Mar 2010
Topic revision: r5 - 03 Mar 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback